### v1

@@ 3.4 Robustness of the control circle
An exciting thought experiment is discussed in [4]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $[Na^{+}]_{o}$, the outside or extracellular sodium concentration, is $145\ mM$, and $[Na^{+}]_{i}$ is 12 mM. And suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential sets up to $66.6\ mV$. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In lack of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law, alone, does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates its electric potential, which is a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electric potential. However, the thermal driving force exists, and the started ion transport quickly builds the electric layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage remains the same, and the voltage defined in this way adjusts the concentrations appropriately. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form a membrane, establishing cellular electricity. As Fig. 6 depicts, the concentrations may adapt to the living conditions (whether living in see-water), but the general operating principles remained the same. The see-water environment enables using thinner membrane and lower membrane voltage; however, the same electric field is is used in all cases.
@@
3.5 Leakage current
Although through $\Delta z$, the ion’s water-related behavior can influence the surface charge density, one can draw the conclusion that when only one type of ion is solved in the liquid, approximately the same concentration difference between the segments produces approximately the same "thermodynamic electric field" that can counterbalance the electric field of the biological condenser. One can notice that in the classic measurement by Hodgkin-Huxley, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) are 380 and 390, respectively, while the sum concentrations of the $Cl^{-}$ and $Ca^{2+}$ are 450 and 460; practically all equal, as our analysis derived it must be. The deviations are well within the uncertainty of the measured values.
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage is generated. For this reason, they assumed that a permanent leakage current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a $42.5\ mV$ voltage [3]. In HH’s picture, the question what ion constitutes the current remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leakage current equals the potassium current. That is, at rest, two currents flow (without driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leak^{+/-}$ ions (??) (depending on the direction, it could be positive or negative), and neither of them changes the concentration neither on the departure nor on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law.
The idea of leakage current was wrong: the voltage HH measured [3] is correct, but – as our model explains – it was generated by charge separation rather than a current through a distributed resistor. It is permanently present while ionic segments are on its sides, even when no current flows. Measuring the energy consumption of neuronal operation [13] confirms that no resting current, in the sense as HH used the notion, exists. The existence of leakage current would be against the overall metabolic efficiency of neuronal operation.
3.6 Heat emission/absorption
The lack of resting current also solves the long-standing mystery of "reversible heat release during the action potential", which is inconsistent with the Hodgkin–Huxley model. In the framework of that theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}*R$) the ion currents flow. As Hodgkin wrote [32]: "Hill and his colleagues found [33] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation."
The issue here is that the dissipation was attributed to the "leakage current." (Although, due to the different conduction mechanisms, an ion current is not dissipated in the way electron energy is dissipated in solids.) However, as our explanation shows, the voltage generated across the membrane is due to charging rather than a permanently flowing current, so no dissipation was present given that no resting current was present. Although the energy to build a potential on the membrane and to maintain it needs biological energy consumption, the process of generating an action potential is, in principle, reversible: ions flow into the layer on the membrane and flow out through the AIS.
According to that, the "temperature of individual particles" with the mean kinetic energy $E$ is $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is the temperature is directly proportional to the kinetic energy. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electric potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature. In the second phase of the AP, those more energetic ions that provide excess voltage above the resting potential leave the membrane through the AIS, so the temperature decreases in the second phase. The heat liberation and absorption are a direct experimental evidence of the "slow current" and the mechanisms described above.
@@ 3.7 Goldman-Hodgkin-Katz potential
In our clear physical picture, the thermodynamical forces on one side of the membrane and on the other side are summed and balance each other. As we emphasized, the concentration gradients are ion specific; furthermore, that the Nernst equation comprises a per-ion indefinite constant. The absolute concentration alone has no meaning for a thermodynamical gradient; furthermore, one cannot find a physical reason why concentration on one side should be intermixed with concentrations on the other side. Combining non-gradient type meaningless concentrations in an expression used for deriving electric gradient is not more than number magic; it has no physical background. Even if we intermix arbitrarily chosen ions from both sides and multiply by a logarithm base. The idea itself is a nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability nor with ion absorption [18]. The causality is reversed: the potential is a static notion and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of forming the gradients, mobility and permeability play a role but not in determining the concentrations or the resting potential.
The $Ca^{2+}$ ions do not fit into the GHK picture. However (see our Table 1), it provides a thermodynamic contribution in the same order of magnitude as the charge separation by the membrane. If we add the potential due to $Ca^{2+}$, we arrive at the correct result that the total balance of the thermodynamic contributions equals zero, that is, the resting potential is zero. The electric potential contribution is an additional term, resulting in the resting potential.
It was already stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions adsorption on the membrane surface generates the membrane potential", for a review see [2]. In the light of our findings, we must reconsider a few key concepts. We derived (see Eq.(19)) that the gap voltage depends on the thickness of the membrane and the overall concentration. Furthermore, we demonstrated that the gap voltage is equivalent to the resting potential (correctly: potential difference) and is independent of ion mobility or ion composition in the solution.
When adding a new ion to the solution, the sum concentration increases, and so the electric force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$). No linear summing or similar assumptions can be made. The potential is described by coupled equations as discussed in connection with Eq.(34). As we discussed in connection with Eq.(2), the integral form of the Nernst equation implies an undetermined additive constant. To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is changing the base of the logarithm used in a calculation to match the experimental value.
3.8 Reversal potential
From our picture, it is clear that the reversal potential is reached when the potential on the inner segment is lowered or lifted in such a way that the orange sloping potential in Fig. 7 turns into horizontal. To reach it (provided that the potential on the other side is kept fixed), the concentrations on the invasion site must be changed. The invasions resulting in the exact (but opposite) change of potentials from the two sides leads to different concentrations. The invasion acts as changing the potential at the corresponding surface side of the membrane. Given that the difference between the two potentials on the two surfaces defines the driving forces, the concentration on the other side is also affected. That is, the reversal potential exists, but its value depends on the context, see the discussion about invasions.
4 Moving in the membrane’s potential
In Fig. 4, an ion channel is depicted in the middle of the figure with a diameter of about $3\ nm$ for visibility. Furthermore, we assume that the ion’s size and, correspondingly, the thickness of the atomic layer in the electrolyte on the surface of the membrane is about $0.1\ nm$ (although with their electric double layers, they can also reach the size of $1\ nm$). For comparison, recall that the size of the tip of the clamp pipette is in the range of $1,000\ nm$ and the size of the soma in the range of $10,000\ nm$.
4.1 Passing through ion channels
Figure 4 also hints at how ions can move in the proximity of the membrane. On the left side, the electric field increases toward the membrane, so the ions cannot move in that direction without an initial force being applied. The case is the same on the right side, because the charge is the opposite. An ion must gain energy to get closer to the membrane, which means, at the same time, higher potential energy. This situation explains why the ions do not diffuse into the bulk region. At their present positions, the counterforces by the membrane (and the dipole layer on it) balance the attraction from the opposite side.
Notice that the resulting electric field cannot accelerate an ion except when the ion is at the beginning of the open ion channel across the membrane. Here, the mechanical counterforce is missing, and the field drops to the value generated by the condenser inside the plates. Here, a vast electric field accelerates the ions to a high speed (although a force needed to move against the viscosity decreases the electrostatic acceleration). The empty red circles show an ion that traverses through the channel. The figure shows their electric fields on the two sides, at the beginning and end of the channel. (On the right side, the electric field is negative, but so is the charge of the ions.)
4.2 Moving in the polarized electrolyte
The electric acceleration sharply decreases after leaving the ion channel (see the right side). As shown by the difference of the vertical positions of the filled red and green balls, the difference in the electric field values is huge across the membrane ($E_{accelerate}$) and considerable inside the dielectric electrolyte ($E_{assist}$) as shown by the difference in the electric field values of the green balls. The field still accelerates, but significantly less intensely than the one in the gap. Suppose we assume that the ion quickly accelerates to its Stokes-Einstein speed. In that case, the travel speeds in those space regions are proportional to the difference of the fields in the shown positions and the reciprocals of their travel times.
The passed ions decrease the concentration and the electric field on the high-concentration side, and, concersely, increase them on the low-concentration side. Similarly, the local gradients decrease and increase. At the exit, the field continues to accelerate the ion. However, the increasing potential and concentration increasingly decelerate the passing ions, so they quickly brake them to the assisted speed. Then, the effect of the electric field cancels, and the ions will move with their diffusion speed ("the ion stops"). The local potentials, connected to the operation of the ion channels, act as an "electric lock".
4.3 Ion selectivity
As Eq.(6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamical forces are equal. The $E_{thermal}^{C_{k}}(d)$electric field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions are different. While one ion experiences a resulting force and moves to the other segment, the other remains resting. Of course, the ions moving across the segments change the overall concentrations in both segments (and so the electric field contribution), indirectly affecting the balanced state of the other ions. When a mixture of ions is present in the volume, all ions feel a different driving force, and the solution as a whole will be balanced when the resulting forces for all chemical ions are balanced. The permanently open ion channels only passively participate in the process. The driving forces of the ions are different, and they produce, depending on the actual concentrations, the selectivity of the channels.
When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential. All other ions, including $K^{+}$, feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave controversially during the transient period. The actual voltage gradient may temporarily direct the chemical gradient in opposite directions. Phenomena such as ’over-relaxation’, in this sense, similar to ’hyperpolarization’, may happen.
However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or a few atomic layers on top of each other, depending on the concentration and voltage in the bulk on the two sides. (The voltage gradient diagram line validates the plane crossing the membrane and the ion channel.)
Our results align with the observation (see caption of 11.22 in [26]). The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a large change in the membrane potential. Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions.
@@ 5 Summary
By introducing theoretically that biological currents are slow we underpinned the experimental observation by Hodgkin and Huxley [34] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Recognizing that biological systems use charged ions instead of electrons solves some old mysteries, and defied old fallacies.
@@
By understanding that the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions) defines the value of the resting potential we defied the decades-old fallacy that some mystic concerted operation of ion channels and some linear combination of purely-defined permeability values define the resting potential of neurons. We calculated by using the principles of classic electricity that neurons handle charge as electricity describes (electrolyte) condensers. We introduced a resting and an excited state and established that that the neuron works as a simple control circuit where the set-point is defined by the construction and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat absorption" by neurons, the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH guessed in the physical picture behind their empirical model and mathematical equations, and introduced the correct physical model by providing also its detailed mathematical description.
@@
We could theoretically underpin (see Fig. 6) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH, although not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation on the membrane, as well as polarization in the electrolyte surrounding the membrane. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Considering the presence of electrolytes on the two sides of the membrane, a correct value for the resting potential can be derived numerically from the first principles. Those layers have a decisive impact on the neuron’s electrical operation. The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [7] of science must be considered.
We have shown that the neuron is a well-designed, simple control circuit with separate ground and excited state mechanisms. The geometry and the overall concentration of the cellular fluids determine the nominal values of the electric and thermal gradients. The ion currents through the permanently open "resting" channels and the gated "transient" ion channels, respectively, deliver ions, which means they affect the gradients until the equilibrium state is kept or restored.

### v2

@@ 5.2 Robustness of the control circle
An exciting thought experiment is discussed in [5]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $Na^{+}]_{o}$, the outside or extracellular sodium concentration, is 145mM145\ mM, and $Na^{+}]_{i}$ is 12 mM. And suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential is set up to 66.6mV66.6\ mV. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In the absence of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law alone does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates its electric potential, which is a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electric potential. However, the thermal driving force exists, and the started ion transport quickly builds the electric layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage does not depend on its size, so it remains constant, and the voltage defined in this way adjusts the concentrations accordingly. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form an imperfect membrane with holes, establishing cellular electricity. As Fig. 9 depicts, the concentrations may adapt to the living conditions (whether living in seawater), but the general operating principles remain the same. The seawater environment enables the use of thinner membranes and lower membrane voltages; however, the same electric field is used in all cases.
@@ 5.3 Goldman-Hodgkin-Katz potential
It was already stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions adsorption on the membrane surface generates the membrane potential", for a review see [2]. The idea itself is nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability nor with ion absorption [20]. Furthermore, there is no idea in GHK about whether the ’setpoint’ (why that specific concentration or potential difference) is present and why the same potential is reset after rough perturbations such as issuing an AP or replicating a cell. The causality is reversed: the potential is a static concept and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of forming the gradients, mobility and permeability play a role, but not in determining the concentrations or the resting potential.
In our clear physical picture, the thermodynamic forces on the one side and on the other side of the membrane are summed. They counterbalance each other; furthermore, jointly the effect of the neuronal condenser, see Fig. 9. As we emphasized, the concentration gradients are ion-specific. Furthermore, as we discussed in connection with Eq.(2), the Nernst equation comprises a per-ion indefinite constant (a potential difference). To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is adding absolute concentrations on the different sides of the membrane, or changing the base of the logarithm used in a calculation to match the experimental value. It is not more than number magic. The potential is described by coupled equations as discussed in connection with Eq. (39).
The $Ca^{2+}$ ions do not fit into the GHK picture. One of the reasons why GHK cannot be good is that $Ca^{2+}]$ is omitted. Biologically, it is hard to believe that $Ca^{2+}$ does not participate in the game of life (especially since biology sees the need for $Ca^{2+}$ pumps). Physically, one concentration alone on one side of the membrane cannot keep balance; as the different ion exchange processes, Table 1 and Fig. 9 show. When adding a new ion to the solution, the sum concentration increases, and so the electrical force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$).
@@ 6 Summary
We provided a holistic, multi-level view of neuronal operation and described quantitatively the actual physical processes (instead of biophysics’ fairy tales about hypothesized lipid mechanisms); this way, we provided the promised ’new understanding’ (and the need for rewriting books in neuroscience). The classical neuron model, despite its ambition to be an electrical model, contradicts fundamental laws of electricity and conservation laws of science; furthermore, it cannot interpret other (thermodynamic, mechanical, optical) disciplinary phenomena. The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We proved that no compromise is needed: the pressure and the voltage observe the same physical process by different means. Neuroscience forgot that the neural current consists of ions, and they repel each other, in a confined volume.
In simple words, the neuron combines different disciplines; it works like a two-tact biological internal combustion engine, and can be described as a special Carnot-type thermodynamic engine. The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a vast pressure impulse, due to the electric repulsion of ions. The neuron’s volume remains practically unchanged, and the pressure wave moves the electrolyte (containing dissociated ions) out of the neuron through the only available output channel: the axon. In the first stroke, the pressure impulse invests energy into compressing the elastic membrane and starting a compression wave (soliton). In the second stroke, the elastic membrane pumps out electrolyte through the AIS, which "resists", and this way generates a potential wave, known as AP. The compression produces heat, observable as a temperature increase, and the expansion consumes heat, observable as a temperature decrease; these are well-understood processes in thermodynamics. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). In the discipline of electricity (when using the serial $RC$ model), it can be interpreted as a capacitive current, which in physiology is observed as hyperpolarization that inverts the polarity of the output voltage. The ions represent both mass and charge simultaneously, and correspondingly, in addition to the electric potential wave, mechanical, optical, and other feature changes can be observed. In different disciplines, pressure and voltage represent the same phenomenon, using concepts from various disciplines.
Research already provided the fundamental principles of a correct new ‘net electric’ neuron model, the laws of motion for describing biology, and provided initial steps toward a correct cross-disciplinary model (promising a unified and scientifically complete electric/thermodynamic model); all these open new perspectives for making exploratory experimental and theoretical research. The perspective includes an understanding of the actual thermodynamic and mechanical behavior of cells (and possibly, life) and the relationship between neural information and its carriers. The fact that the neuron receives through its synapses less than 3% of the energy it consumes during its AP cycle underscores that a neuron is a “summing amplifier” rather than a “repeater”. The time window is very narrow compared to the length of the operating cycle, so the neuron receives inputs only from a small subset of neurons connected to its synapses (the late input signals are gated, as Fig. 12 shows).
@@
We have shown that the neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms and established that the neuron works as a simple control circuit where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentration of the cellular fluids define the nominal values of the electric and thermal gradients. Our results are in line with the conclusions "It is shown theoretically that the membrane potential enhancement becomes greater with decreasing the electrolyte concentration and pore radius" [19]. The ion currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By understanding that the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions) defines the value of the resting potential, we defied the decades-old fallacy that some mystic concerted operation of ion channels and some linear combination of purely-defined permeability values (the GHK equation) define the resting potential of neurons. By using the principles of classic electricity, we calculated, that neurons handle charge as electricity describes (electrolyte) condensers. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat absorption" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description.
@@
We could theoretically underpin (see Fig. 9) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[1]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation on the membrane, as well as polarization in the electrolyte surrounding the membrane. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Considering the presence of electrolytes on the two sides of the membrane, a correct value for the resting potential can be derived numerically from the first principles. Those layers have a decisive impact on the neuron’s electrical operation. The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [8] of science must be considered.
By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [65] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Recognizing explicitly that biological systems use slow currents of charged ions instead of electrons solved some old mysteries and defied old fallacies.
Appendix A Speculations
By using the value of the force acting on a unit charge in the field across the membrane is
$F_{Na^{+}}=10^{7}*1.60217663*10^{-19}[V][C]=1.6*10^{-12}\ [N]$ (48)
we can estimate how the pressure of the neural cell increases due to the rush-in changes at the beginning of the AP. As evidence shows, the local potential at the internal surface of the membrane increases by $\Delta U=100\ [mV]$. This increase means a change in the force acting on an ion (see Eq. (48)) by $1.6*10^{-12}\ [N]$. When we assume $10^{7}$ rush-in ions and unchanged cell size, the total force acting on the membrane increases by $1.6*10^{-5}\ [N]$. This change in force means on the neuron’s $8*10^{-9}\ [m^{2}]$ surface (see [30], page 12) a pressure change
$\Delta P=\frac{1.6*10^{-5}\ [N]}{8*10^{-9}\ [m^{2}]}=2*10^{3}\quad\biggl[\frac{N}{m^{2}}\biggr]$ (49)
[…]
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed ([49], Fig. 3 quantitatively underpins the hypothesis) both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [49] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$ the speed at the exit of the ion channels can be estimated as $10^{4}~\frac{m}{s}$.
We can also estimate the work done when the rush-in charge extends the neuron, as the $\Delta P\times\Delta V$. We consider that $\Delta P$ remains constant, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $\Delta P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [70], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$. We can also estimate the energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [14]. To estimate the time of changing the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave) and it can be calculated as $F=m\frac{\Delta v}{\Delta t}$. Given that Eq. (48) provides the force and one must assume that the ion loses its $2*10^{5}~\frac{m}{s}$ speed, one arrives that it generates the shock wave in $\Delta t=5*10^{-8}~seconds$.
The consequences of the changes in the electric charge are large enough to explain why pressure wave and other mechanical changes [37] also start at the beginning of the AP, and other (such as optical, density) changes are accompanied by it. Nature invests energy also in the conventional way, a term $V\Delta P$, into the thermodynamics of neural operation; not only in the form of storing energy in the changed electric field. Thermodynamics and electricity, not to mention elasticity, must not be separated when discussing the neuronal energy business.
In line with the experimentally measured energy consumption and our calculation in section 4.5, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. That is, practically all the energy is invested as elastic energy, but the modulus is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in proximity to the membrane, but it creates a secondary, vast mechanical shock wave. By considering the electric repulsion force between the particles in the slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion consumes is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electric potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature (the temperature is more that two orders of magnitude higher, but the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect temperature change up to the range of fragments of millikelvins, depending on the measuring conditions). In the second phase of the AP, those more energetic ions that provide excess voltage (and so: excess local potential) above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value; another effect that can contribute to the general heat production and adsorption discussed above. In the original publication about solitons (essentially pressure waves) [78], a lot of measured and theoretical data are discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change; furthermore, that the electric energy is more than an order of magnitude smaller than the elastic energy (called soliton energy there).